# references.bib

### v1

@@ 4.3.3 Building a structure
4.4 Direct Stiffness method
@@
4.4.1 Generalized Euler-Bernoulli beam elements
In this work, we use generalized Euler-Bernoulli beam elements to model the beams and levers in a Totimorphic lattice. The stiffness matrix $\boldsymbol{K}_{ij}$ of the generalized beam element is obtained by combining the stiffness matrices of rod elements, i.e., elements that behave like springs,
@@ 4.4.1 Generalized Euler-Bernoulli beam elements
and Euler-Bernoulli beam elements that model bending [ochsner2018finite]
@@ 4.4.2 Compression experiment
To implement the compression experiment, we first set the displacement of points that are part of the top surface of the lattice to a small non-zero value. As boundary conditions, we set the displacement of all points in the bottom surface to zero (otherwise, the lattice would start moving downwards). Using Eq. 43 (i.e. matrix multiplication), we calculate the resulting forces for all other points due to these forced displacements – from which we subsequently calculate the resulting displacements of all remaining points in the lattice that are not part of the top or bottom surface by solving a system of linear equations. We repeat this process several times (with updated global stiffness matrix $\boldsymbol{G}$) to obtain macroscopic deformations. For a detailed description, see [dold2023differentiable].
@@ 4.4.3 Poisson’s ratio
where $\mathcal{LS}$ ($\mathcal{RS}$) is a set containing the indices of the points on the left (right) surface of the lattice. $|\mathcal{RS}|$ denotes the number of elements in $\mathcal{RS}$. The Poisson’s ratio is obtained using linear regression
@@
4.5 Ray-based optics
In the telescope mirror application (Section 2.4), for simplicity, we only simulate light rays getting reflected at the mid-point (i.e. the mid-point of the lever) of each unit cell. Thus, each unit cell acts as a tiny mirror reflecting a single beam of light. We assume light coming from a source that is far away, such that the light rays hitting the different mirror elements are parallel to each other.
@@ 4.5.1 Reflections
$\boldsymbol{n}_{ij}$ $=\frac{\left(\boldsymbol{r}^{\mathrm{B}}_{ij}-\boldsymbol{r}^{\mathrm{A}}_{ij}\right)\times\left(\boldsymbol{r}^{\mathrm{C}}_{ij}-\boldsymbol{r}^{\mathrm{A}}_{ij}\right)}{\|\left(\boldsymbol{r}^{\mathrm{B}}_{ij}-\boldsymbol{r}^{\mathrm{A}}_{ij}\right)\times\left(\boldsymbol{r}^{\mathrm{C}}_{ij}-\boldsymbol{r}^{\mathrm{A}}_{ij}\right)\|_{2}}\,,$ (48)
$\boldsymbol{n}^{\prime}_{ij}$ $=\frac{\left(\boldsymbol{r}^{\mathrm{B^{\prime}}}_{ij}-\boldsymbol{r}^{\mathrm{C}}_{ij}\right)\times\left(\boldsymbol{r}^{\mathrm{P^{\prime}}}_{ij}-\boldsymbol{r}^{\mathrm{C}}_{ij}\right)}{\|\left(\boldsymbol{r}^{\mathrm{B^{\prime}}}_{ij}-\boldsymbol{r}^{\mathrm{C}}_{ij}\right)\times\left(\boldsymbol{r}^{\mathrm{P^{\prime}}}_{ij}-\boldsymbol{r}^{\mathrm{C}}_{ij}\right)\|_{2}}\,,$ (49)
@@ 4.5.2 Smallest distance to the target focal point
and optimise the Totimorphic lattice by updating its parameters via gradient descent $\nabla_{\mathcal{P}}C(\mathcal{P},\boldsymbol{f}^{\mathrm{tgt}})$.
@@ 4.5.3 Damaging mirror elements
To emulate a damaged mirror element, we add a random perturbation to the direction of the reflected light. For instance, assuming the bottom unit cell of element $(ij)$ was damaged, we first sample a random deflection $\boldsymbol{l}_{\sigma}\sim\mathcal{N}(0,\sigma^{2})$ from a Gaussian distribution (with mean 00 and variance $\sigma^{2}$) and always add it to the reflected light direction:
@@ 4.5.4 Imaging
For a $2S\times 2S$ pixel image of Pluto, the light ray coming from the pixel in row $i$ and column $j$ (counted from the bottom left corner of the image) has the an unnormalized incidence vector
@@
and the normalized one $\boldsymbol{I}_{ij}=\frac{\tilde{\boldsymbol{I}}_{ij}}{\|\tilde{\boldsymbol{I}}_{ij}\|_{2}}$. To get an image, we reflect the light rays coming from every pixel in the image at every Totimorphic mirror element and check where on the imaging plane the reflections hit. For instance, if the imaging plane is at distance $d_{z}$, the light ray $\boldsymbol{I}_{kl}$ reflected at the bottom unit cell of element $(ij)$ hits the plane at $\boldsymbol{m}_{ij}+\lambda\cdot\boldsymbol{l}_{ij}$ with $\lambda=\left(d_{z}-m_{ij,z}\right)/l_{ij,z}$ ($z$ denoting the $z$-component of the vectors). For visualisation, we plot all reflected points in a two-dimensional image with point size $s=2$ (in matplotlib) and an alpha (opacity) of $0.05$, so that colours merge when multiple points overlap.
Contributions
Initial idea is by AT and DD. Theory and model implementation is by DD, based on initial work by AT. All authors contributed to the design of the study and the performed experiments. DD and NR implemented code for the proof of concept experiments. DD wrote the first draft of the paper. All authors read and reviewed the final paper.
Acknowledgments
We would like to thank Derek Aranguren van Egmond, Michael Mallon, and Max Bannach for helpful and inspiring discussions, and our colleagues at ESA’s Advanced Concepts Team for their ongoing support. AT, NR, and DD acknowledge support through ESA’s young graduate trainee and fellowship programs. DD further acknowledges support through Horizon Europe’s Marie Skłodowska-Curie Actions (Project 101103062 — BASE).
[…]
Supplemental Information
S1 Proofs for two-dimensional case
S1.1 Finding $\mathrm{P^{\prime}}$
$\|\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{P^{\prime}}}\|^{2}$ $=\left(\frac{l_{\text{b}}}{2}\right)^{2}\,,$ (S1a)
$\|\boldsymbol{r}^{\mathrm{C}}-\boldsymbol{r}^{\mathrm{P^{\prime}}}\|^{2}$ $=\left(\frac{l_{\text{b}}}{2}\right)^{2}\,.$ (S1b)
Expanding both expressions, subtracting Eq. S1b from Eq. S1a and dividing the resulting equation by $2$, we obtain:
$\frac{1}{2}\left(\|\boldsymbol{r}^{\mathrm{A^{\prime}}}\|^{2}-\|\boldsymbol{r}^{\mathrm{C}}\|^{2}\right)-r^{\mathrm{P^{\prime}}}_{y}\left(r^{\mathrm{A^{\prime}}}_{y}-r^{\mathrm{C}}_{y}\right)-r^{\mathrm{P^{\prime}}}_{x}\left(r^{\mathrm{A^{\prime}}}_{x}-r^{\mathrm{C}}_{x}\right)=0\,.$ (S2)
Using the notation $\boldsymbol{s}^{-}=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{C}}$, $\boldsymbol{s}^{+}=\boldsymbol{r}^{\mathrm{A^{\prime}}}+\boldsymbol{r}^{\mathrm{C}}$ and $\delta=\frac{1}{2}\left(\|\boldsymbol{r}^{\mathrm{A^{\prime}}}\|^{2}-\|\boldsymbol{r}^{\mathrm{C}}\|^{2}\right)$ as in the main text, this yields Eq. 24. Eq. 25 is then obtained by inserting this solution into either Eq. S1a or Eq. S1b and solving for $r^{\mathrm{P^{\prime}}}_{y}$. This results in a quadratic equation, possessing two solutions
$r^{\mathrm{P^{\prime}}}_{y}=\frac{1}{2}\left(s^{+}_{y}\pm s^{-}_{x}\sqrt{\left(\frac{l_{\text{b}}}{\|\boldsymbol{s}^{-}\|}\right)^{2}-1}\right)\,.$ (S3)
S1.2 Condition for existing solution
For a solution of $\boldsymbol{r}^{\mathrm{P^{\prime}}}$ to exist, Eq. 25 has to be a real number, meaning that the argument of the square root has to be larger or equal to zero. This immediately yields Eq. 26.
S1.3 Deriving allowed range for $\theta$
First, we write out $\boldsymbol{r}^{\mathrm{C}}$ in $\boldsymbol{s}^{-}$:
$\boldsymbol{s}^{-}$ $=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{C}}\,,$ (S4a)
$=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{A}}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi+\theta}\,,$ (S4b)
$=\boldsymbol{k}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi+\theta}\,,$ (S4c)
with $\boldsymbol{k}=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{A}}-\frac{l_{\text{b}}}{2}\boldsymbol{e}^{\varphi}$. This way, we split up $\boldsymbol{s}^{-}$ into two terms – one depending on $\theta$, and one that is independent of $\theta$, summarizing the geometric constraints introduced by the previous unit cell. By representing $\boldsymbol{k}$ in polar coordinates, we arrive at
$\boldsymbol{s}^{-}=\|\boldsymbol{k}\|\boldsymbol{e}^{\Phi}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi+\theta}\,,$ (S5)
with $\Phi=E\left(\boldsymbol{k}\right)$, which simplifies the following derivation of the range of $\theta$. To guarantee that a solution for point $\mathrm{P^{\prime}}$ exists, Eq. 26 has to be fulfilled. By squaring Eq. 26 and using Eq. S5, we get
$\left\|\|\boldsymbol{k}\|\boldsymbol{e}^{\Phi}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi+\theta}\right\|^{2}\leq l_{\text{b}}^{2}\hskip 11.38109pt$ $\Leftrightarrow\hskip 11.38109pt\|\boldsymbol{k}\|^{2}-l_{\text{b}}\|\boldsymbol{k}\|\boldsymbol{e}^{\Phi}\cdot\boldsymbol{e}^{\varphi+\theta}+\frac{1}{4}l_{\text{b}}^{2}\leq l_{\text{b}}^{2}\,,$ (S6a)
$\Leftrightarrow\hskip 11.38109pt\boldsymbol{e}^{\Phi}\cdot\boldsymbol{e}^{\varphi+\theta}\geq\frac{\|\boldsymbol{k}\|}{l_{\text{b}}}-\frac{3}{4}\frac{l_{\text{b}}}{\|\boldsymbol{k}\|}\,,$ (S6b)
$\Leftrightarrow\hskip 11.38109pt\cos\left(\Phi-\varphi-\theta\right)\geq\frac{\|\boldsymbol{k}\|}{l_{\text{b}}}-\frac{3}{4}\frac{l_{\text{b}}}{\|\boldsymbol{k}\|}\,,$ (S6c)
$\Leftrightarrow\hskip 11.38109pt\cos\left(\Delta-\theta\right)\geq C\,,$ (S6d)
±$\pm\left(\Delta-\theta\right)\leq\arccos{C}\,,$ (S7)
from which we get the upper (choosing --) and lower (choosing $+$) bound of $\theta$ shown in Eq. 29.
S2 Proofs for three-dimensional case
S2.1 Calculating $\phi$ and $\Theta$
Expanding Eq. 34, we get
$\boldsymbol{s}^{-}\cdot\boldsymbol{\xi}^{\phi,\Theta}=-\frac{\|\boldsymbol{s}^{-}\|^{2}}{l_{\text{b}}}\,.$ (S8)
$\boldsymbol{\xi}^{\alpha,\beta}\cdot\boldsymbol{\xi}^{\phi,\Theta}=-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\,,$ (S9)
or, using Eq. 16,
$\boldsymbol{\xi}^{\alpha,\beta}\cdot\boldsymbol{\xi}^{\phi,\Theta}=\cos\left(\alpha-\phi\right)\sin{\beta}\sin{\Theta}+\cos{\beta}\cos{\Theta}=-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\,,$ (S10)
from which we get the solution [^10] The arccos can have two signs, so as in the two-dimensional case, two solutions exist. We found that when starting with the flat surface configuration, a minus sign has to be chosen. However, different from the two-dimensional case, the alternative solution ($+$ instead of --) also represents a valid Totimorphic configuration, although the two obtained configurations are far apart. for $\phi$ depending on $\Theta$:
$\phi=\alpha-\arccos\left(\mathcal{C}\right)\hskip 11.38109pt\text{with}\hskip 11.38109pt\mathcal{C}=-\frac{\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}+\cos{\beta}\cos{\Theta}}{\sin{\beta}\sin{\Theta}}\,.$ (S11)
In the case $\sin{\beta}=0$, $\phi$ can be freely chosen and $\Theta$ has a unique solution $\Theta=\arccos\left(-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}\cos{\beta}}\right)$. In this case, the spring vector $\boldsymbol{s}^{-}$ is parallel to the z-axis. Any solution given by $\Theta$ will now form a triangle, from $\mathrm{A^{\prime}}$ to $\mathrm{P^{\prime}}$, from $\mathrm{P^{\prime}}$ to $\mathrm{C}$, and from $\mathrm{C}$ back to $\mathrm{A^{\prime}}$ (which is parallel to the z-axis). Changing $\phi$ will now rotate this triangle around the z-axis without distorting it, hence there is an infinite amount of valid configurations that lie on a circle around the z-axis.
$-1\leq\mathcal{C}\leq 1\,.$ (S12)
$-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\geq\cos{\beta}\cos{\Theta}-\sin{\beta}\sin{\Theta}\,.$ (S13)
$-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\geq\cos\left(2\pi-\beta-\Theta\right)\,.$ (S14)
which yields
$\arccos\left(-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\right)-\beta\leq\Theta\leq 2\pi-\beta-\arccos\left(-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\right)\,.$ (S15)
$\beta-\arccos\left(-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\right)\leq\Theta\leq\beta+\arccos\left(-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\right)\,.$ (S16)
S2.2 Deriving allowed range of $\theta_{z}$
Finally, we can calculate an allowed range for $\theta_{z}$ (as in the 2D case). This can be derived by requiring that the $\arccos$ in Eqs. S15 and S16 has a valid solution, i.e.,
$\|\boldsymbol{s}^{-}\|^{2}\leq l_{\text{b}}^{2}\,.$ (S17)
First, we rewrite $\boldsymbol{s}^{-}$ by expanding it:
$\boldsymbol{s}^{-}=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{C}}=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{P}}-\frac{l_{\text{b}}}{2}\cdot R_{y}(\varphi_{y})R_{z}(\varphi_{z})R_{x}(\theta_{x})R_{z}(\theta_{z})\mathbb{1}_{x}$ (S18)
$R_{x}(-\theta_{x})R_{z}(-\varphi_{z})R_{y}(-\varphi_{y})\boldsymbol{s}^{-}=\boldsymbol{k}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\xi}^{\theta_{z},\frac{\pi}{2}}\,,$ (S19)
with $\boldsymbol{k}=R_{x}(-\theta_{x})R_{z}(-\varphi_{z})R_{y}(-\varphi_{y})\cdot\left(\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{P}}\right)$. Since rotations leave the length of a vector unchanged, we can thus rewrite condition Eq. S17 as
$\|\boldsymbol{k}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\xi}^{\theta_{z},\frac{\pi}{2}}\|^{2}\leq l_{\text{b}}^{2}\,.$ (S20)
We solve this the same way as in Section S1.3, resulting in
$\boldsymbol{\xi}^{\theta_{k},\phi_{k}}\cdot\boldsymbol{\xi}^{\theta_{z},\frac{\pi}{2}}\geq\frac{\|\boldsymbol{k}\|}{l_{\text{b}}}-\frac{3}{4}\frac{l_{\text{b}}}{\|\boldsymbol{k}\|}\,,$ (S21)
[…]
from which the final interval is derived
$\theta_{k}-\arccos{C}\leq\theta_{z}\leq\theta_{k}+\arccos{C}\,.$ (S23)
S3 Simulation details
S3.1 Inverse lattice design
@@ Acknowledgments
S3.2 Totimorphic telescope
We trained the lattice only on the task of focusing light from one direction (i.e., orthogonal to the initially flat mirror surface) into a single point (focal point on the central axis orthogonal to the initially flat mirror surface) located at $(\frac{D}{2},\frac{D}{2},n\cdot D$, where $D=6$m is the diameter (or side length) of the Totimorphic lattice and $n$ an integer that we choose during training (e.g., we start with $20$). For training configurations at different focal lengths, we used the Adam optimiser with a learning rate of $10^{-3}$, which is reduced to $10^{-4}$ if the root mean square loss goes below $5\cdot 10^{-3}$. We further always add a second cost term (with prefactor of $0.1$) which has the objective of keeping the central node (point $\mathrm{B^{\prime}}$ of the central elementary cell) of the lattice fixed in space. This is done to avoid that the lattice moves around in space during optimisation.

### v2

@@ 4.3.5 Building a three-dimensional Totimorphic structure
Depending on where in the lattice an elementary cell is positioned, it is characterised by the following parameters (we drop indices here for clarity, but all parameters are different):
@@
4.4 Methodology for the first proof of concept: direct stiffness method
@@
4.4.1 Generalised Euler-Bernoulli beam elements
In this work, we use generalised Euler-Bernoulli beam elements to model the beams and levers in a Totimorphic lattice. The stiffness matrix $\boldsymbol{K}_{ij}$ of the generalised beam element is obtained by combining the stiffness matrices of rod elements, i.e., elements that behave like springs,
@@ 4.4.1 Generalised Euler-Bernoulli beam elements
and Euler-Bernoulli beam elements that model bending [17]
@@ 4.4.2 Compression experiment
To implement the compression experiment, we first set the displacement of points that are part of the top surface of the lattice to a small non-zero value. As boundary conditions, we set the displacement of all points in the bottom surface to zero (otherwise, the lattice would start moving downwards). Using Eq. 105 (i.e. matrix multiplication), we calculate the resulting forces for all other points due to these forced displacements – from which we subsequently calculate the resulting displacements of all remaining points in the lattice that are not part of the top or bottom surface by solving a system of linear equations. We repeat this process several times (with updated global stiffness matrix $\boldsymbol{G}$) to obtain macroscopic deformations. For a detailed description, see [4].
@@ 4.4.3 Poisson’s ratio
where $\mathcal{LS}$ ($\mathcal{RS}$) is a set containing the indices of the points on the left (right) surface of the lattice. The outer points are given (left side) by points $\mathrm{A}$ and $\mathrm{A^{\prime}}$ of the first (i.e. outer left) column of elementary cells in the lattice, and (right side) by the points $\mathrm{B}$ and $\mathrm{B^{\prime}}$ of the last column (i.e. outer right) of elementary cells. $|\mathcal{RS}|$ denotes the number of elements in $\mathcal{RS}$. The Poisson’s ratio is obtained using linear regression
@@
The deformation in vertical direction is imposed on the top points, i.e., $\mathrm{A^{\prime}}$ and $\mathrm{B^{\prime}}$ of all elementary cells in the top roow, while the bottom points ($\mathrm{A}$ and $\mathrm{B}$ of all elementary cells in the bottom row) are kept static.
Alternative ways of calculating the Poisson ratio are possible, e.g., by imposing target displacements for the outer nodes in the cost function (e.g., no displacement to reach a Poisson ratio of 0, as done in [4], Fig. 6C). Furthermore, to enforce symmetric behaviour during compression, additional terms can be added to the cost function that punish asymmetric deformations under load.
4.5 Methodology for the second proof of concept: ray-based optics
In the telescope mirror application, for simplicity, we only simulate light rays getting reflected at the mid-point (i.e. the mid-point of the lever) of each unit cell. Thus, each unit cell acts as a tiny mirror reflecting a single beam of light. We assume light coming from a source that is far away, such that the light rays hitting the different mirror elements are parallel to each other.
@@ 4.5.1 Reflections
$\boldsymbol{n}_{ij}$ $=\frac{\left(\boldsymbol{r}^{\mathrm{B}}_{ij}-\boldsymbol{r}^{\mathrm{A}}_{ij}\right)\times\left(\boldsymbol{r}^{\mathrm{C}}_{ij}-\boldsymbol{r}^{\mathrm{A}}_{ij}\right)}{\left\|\left(\boldsymbol{r}^{\mathrm{B}}_{ij}-\boldsymbol{r}^{\mathrm{A}}_{ij}\right)\times\left(\boldsymbol{r}^{\mathrm{C}}_{ij}-\boldsymbol{r}^{\mathrm{A}}_{ij}\right)\right\|_{2}}\,$, (110)
$\boldsymbol{n}^{\prime}_{ij}$ $=\frac{\left(\boldsymbol{r}^{\mathrm{B^{\prime}}}_{ij}-\boldsymbol{r}^{\mathrm{C}}_{ij}\right)\times\left(\boldsymbol{r}^{\mathrm{P^{\prime}}}_{ij}-\boldsymbol{r}^{\mathrm{C}}_{ij}\right)}{\left\|\left(\boldsymbol{r}^{\mathrm{B^{\prime}}}_{ij}-\boldsymbol{r}^{\mathrm{C}}_{ij}\right)\times\left(\boldsymbol{r}^{\mathrm{P^{\prime}}}_{ij}-\boldsymbol{r}^{\mathrm{C}}_{ij}\right)\right\|_{2}}\,$, (111)
@@ 4.5.2 Smallest distance to the target focal point
and optimise the Totimorphic lattice by updating its parameters via gradient descent, $\nabla_{\mathcal{P}}C(\mathcal{P},\boldsymbol{f}^{\mathrm{tgt}})$.
@@ 4.5.3 Damaging mirror elements
To emulate a damaged mirror element, we add a random perturbation to the direction of the reflected light. For instance, assuming the bottom unit cell of element $(ij)$ was damaged, we first sample a random deflection $\boldsymbol{l}_{\sigma}\sim\mathcal{N}(0,\sigma^{2})$ from a Gaussian distribution (with mean 0 and variance $\sigma^{2}$) and always add it to the reflected light direction:
@@ 4.5.4 Imaging
For a $2S\times 2S$ pixel image of Pluto, the light ray coming from the pixel in row $i$ and column $j$ (counted from the bottom left corner of the image) has the non-normalised incidence vector
@@
and the normalised one $\boldsymbol{I}_{ij}=\frac{\tilde{\boldsymbol{I}}_{ij}}{\|\tilde{\boldsymbol{I}}_{ij}\|_{2}}$. To get an image, we reflect the light rays coming from every pixel in the image at every Totimorphic mirror element and check where on the imaging plane the reflections hit. For instance, if the imaging plane is at distance $d_{z}$, the light ray $\boldsymbol{I}_{kl}$ reflected at the bottom unit cell of element $(ij)$ hits the plane at $\boldsymbol{m}_{ij}+\lambda\cdot\boldsymbol{l}_{ij}$ with $\lambda=\left(d_{z}-m_{ij,z}\right)/l_{ij,z}$ ($z$ denoting the $z$-component of the vectors). For visualisation, we plot all reflected points in a two-dimensional image with point size $s=2$ (in matplotlib) and an alpha (opacity) of $0.05$, so that colours merge when multiple points overlap.
4.6 Simulation details: inverse lattice design
@@
4.7 Simulation details: Totimorphic telescope
We trained the lattice only on the task of focusing light from one direction (i.e., orthogonal to the initially flat mirror surface) into a single point (focal point on the central axis orthogonal to the initially flat mirror surface) located at $(\frac{D}{2},\frac{D}{2},n\cdot D$, where D=6D=6m is the diameter (or side length) of the Totimorphic lattice and $n$ an integer that we choose during training (e.g., we start with $20$). For training configurations at different focal lengths, we used the Adam optimiser with a learning rate of $10^{-3}$, which is reduced to $10^{-4}$ if the root mean square loss goes below $5\cdot 10^{-3}$. We further always add a second cost term (with prefactor of $0.1$) which has the objective of keeping the central node (point $\mathrm{B^{\prime}}$ of the central elementary cell) of the lattice fixed in space. This is done to avoid that the lattice moves around in space during optimisation.